40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat May 9 00:15:21 2009, Yoichi, Update, PSL, Laser head temperature oscillation laser.png
    Reply  Sat May 9 02:20:11 2009, Jenne, Update, PSL, Laser head temperature oscillation 
       Reply  Sat May 9 15:19:10 2009, rana, Update, PSL, Laser head temperature oscillation Picture_2.png
          Reply  Mon May 11 11:49:20 2009, steve, Update, PSL, MOPA cooling water lines are backwards htempvar80d.jpg
             Reply  Thu May 21 18:07:32 2009, rana, Update, PSL, Screw on Needle valve loosened Untitled.png
Message ID: 1617     Entry time: Thu May 21 18:07:32 2009     In reply to: 1573
Author: rana 
Type: Update 
Category: PSL 
Subject: Screw on Needle valve loosened 
Alberto and I went in to loosen up the needle valve yesterday around 4:30 PM. The idea was to cut down on
the flow to the NPRO so that the cooling power of the chiller would be used almost entirely on the
amplifier instead of the NPRO block.

The need valve was basically all the way open. The lock nut was screwed in all the way and stuck. By using
pliers and a wrench for the nut, we freed the lock nut. Even so, the screw for the needle valve seemed to
be bad: I think the thread is stripped; it doesn't go down even after several turns. I even tried to squirt
alchohol on it and really press down in the hopes of catching a thread. It may have closed slightly but its
impossible to be sure.

I also increased the NPRO diode current to the max (+0.1 A). This got us a little bit of NPRO power and
I hope some more AMPMON stability. The attached plot shows 4 days of minute trend. If you squint you
might believe that we got some suppression in the HTEMP fluctuations over the last two days.
Attachment 1: Untitled.png  148 kB  Uploaded Thu May 21 19:33:20 2009  | Hide | Hide all
Untitled.png
ELOG V3.1.3-